2x + 3y = 12
4x - 3y = 6
what is the solution and how do i plot 3 points on the graph to make a line?

Answer by rfer(16322)

create an input/output table for each equation, then plot the points
I will get you started.
2x+3y=12
(6,0)
(0,-2)
-----------
4x-3y=6
(1.5,0)
(0,-2)
they intersect at (0,-2)